You have an opportunity to buy a building. You have a tenant lined up that will generate $20,254 per year in cash flows for two years. The rent pays at the beginning of each year. At the end of two years, you anticipate selling the building for $248,962. How much would you be willing to pay for the building if the discount rate is 8%?
Answer #1

Willing to pay = (1 + r) * Annuity * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n

Willing to pay = (1 + 0.08) * 20,254 * [1 - 1 / (1 + 0.08)2] / 0.08 + 248,962 / (1 + 0.08)2

Willing to pay = 1.08 * 20,254 * [1 - 0.857339] / 0.08 + 213,444.7874

Willing to pay = 1.08 * 20,254 * 1.783263 + 213,444.7874

Willing to pay = $252,452.49